The table below shows the income for each person in an economy. Use this information to answer the following questions AssumeQuestion 6 1 pts Assuming we can do perfectly targeted transfers, how much money do we need to reduce poverty to zero?
Answer #1

Answer:

Given, Poverty line Z=6

Hence 3 persons come under the poverty line and their income gap G= (6-3)+(6-4)+(6-5)=6 (as shown by blue color).

Individual Income (Y) Poverty Line (Z) Poor (YZ) Income Gap (G=Y-Z) 3.0 2.0 U AU 6.0 6.0 Total Income Shortfall Average Incom

Hence if there is perfectly targeted transfers, we need 6 ($6 if the currency is dollar) to reduce poverty to zero.